Jane invested some amount at the rate of 12% simple interest and some other amount at 10% simple interest. She received yearly interest of $130. Randy also invested, and received $4 more as interest. How much amount did each of them invest at differen rates?
Answer by checkley71(8403)

I apologize. Randy also invested in the same scheme, but he interchanged the amounts invested, and received $4 more as interest. How much amount did each of them invest at different rates?
.12X+.10Y=130 THE FORMULA FOR JANE'S INVESTMENT.
.10X+.12Y=134 THE FORMULA FOR RANDY'S INVESTMENT.
MULTIPLY JANE'S FORMULA BY -10 & RANDY'S BY 12 THEN ADD THEM.
-1.2X-1Y=-1300
1.2X+1.44Y=1608
------------------------------
.44Y=308
Y=308/.44
Y=$700 INVESTED BY JANE @ 10% & BY RANDY @ 12%. ANSWER.
.12X+.10*700=130
.12X+70=130
.12X=130-70
.12X=60
X=60/.12
X=$500 INVESTED BY JANE @ 12% & BY RANDY @ 10%. ANSWER.
PROOF:
.10*500+.12*700=134
50+84=134
134=134